OpenStudy (herpderp):

Algebra Help Consider the system shown: x - 3y + z = 6 x + 4y - 2z = 9 How many solutions does this system have? Make a conjecture about the minimun number of equations that a linear system in n variables can have when there is exactly one solution.

OpenStudy (jiteshmeghwal9):

since \(\Large{\frac{a_1}{a_2} \ne \frac{b_1}{b_2} \ne \frac{c_1}{c_2}}\) therefore the system possesses a unique solution

OpenStudy (jiteshmeghwal9):

Also to find the solution of a linear system consisting of n variables u need minimum of n equations

OpenStudy (herpderp):

I don't get the first part D:

OpenStudy (wolf1728):

In order to find a solution, you need 1 equation for each unknown. Since there are 3 unknowns you need a third equation.

OpenStudy (danjs):

Underdetermined system...more variables than equations. Can have either infinite number of solutions or zero solutions.

OpenStudy (herpderp):

Yes, I get the part about n equations for n variables

OpenStudy (jiteshmeghwal9):

\[a_1x_1+b_1y_1+c_1z_1=d_1\]\[a_2x_2+b_2y_2+c_2z_2=d_2\] \[\frac{a_1}{a_2} \ne \frac{b_1}{b_2} \ne \frac{c_1}{c_2} \] means that the system has a unique solution

OpenStudy (jiteshmeghwal9):

\[\frac{a_1}{a_2} = \frac{b_1}{b_2} = \frac{c_1}{c_2}\]means that the system possesses infinite number of solutions

OpenStudy (jiteshmeghwal9):

\[\frac{a_1}{a_2}=\frac{b_1}{b_2} \ne \frac{c_1}{c_2}\]means that the system possesses no solutions

OpenStudy (herpderp):

Thanks O_O that's weird

OpenStudy (kevin):

Except if you have 1 equation more. It can be have just one solution.

OpenStudy (herpderp):

What does that mean? @Kevin

OpenStudy (3mar):

By the way, Kevin Even it has one more equation, it is not a condition to has a unique solution, may be none

OpenStudy (kevin):

This is linear equation with 3 variables, you need 3 equations to find each variable values where there is exactly one solution.